# Overcoming the Thermal-Noise Limit of Microwave Measurements by Pre-cooling with an Active Cold Load

### v2

@@ IV Switchover followed by Interrogation
The well-known Landau-Zener (L-Z) modelZener (1932); Glasbrenner and Schleich (2023) provides quantitative insight into the location of this balance. Since, for the relatively high-$Q$ (by room-temperature standards) sapphire-loaded dielectric resonator that we use in our demonstration, the distances in frequency between the operational microwave mode and its neighboring ones correspond to many thousands of cavity linewidths, a switching time that is a tenth or a hundredth of the cavity lifetime easily satisfies adiabaticity according to the L-Z formula (for the probability of a non-adiabatic transition). Note also that even if adiabaticity is compromised by using a very high over-coupling factor (so perturbing the frequencies and shapes of the cavity electromagnetic modes) and/or a too-rapid switchover, the pre-cooling effect is still protected by the fact that those modes closest in frequency to the operational mode, and thus those that stand to get substantially mixed with it (at anti-crossings), will also lie within the bandwidth of the ACL, and thus stand to be cooled by it also (though admittedly their couplings to the ACL and thus their cooling depths will vary). The mixing in of these nearby modes with the operational mode should thus not substantially raise the temperature. In practice, there is a limit to the amount of switchable external coupling (usually in the form of metal stubs or loops) that can be added to a cavity without reducing the intrinsic quality factor (i.e. photon lifetime) of its modes when the external coupling is switched out (disconnected). The frequency shift $\Delta f$ associated with the imperfect switching in and out of an overcoupled ACL, is shown in Fig. 5. If the cavity’s mode structure gets excessively scrambled, more distant “hot” modes beyond the ACL’s cooling bandwidth risk getting mixed into the operational mode at switchover so raising its temperature. We point out here parenthetically that the engineering challenge of achieving a large change in coupling while incurring only a small $\Delta f$ (or at least the avoidance of anticrossings and mode scrambling) draws parallels to the realization of switchable attenuators exhibiting low dispersion, for which proven design strategies/solutions are known. Based on our crude L-Z analysis, there would appear to be at least two decades of headroom for raising $\kappa_{\text{over}}$ to the point of “diminishing returns”, where it is no longer the external coupling but the noise temperature of the room-temperature ACL itself that limits the cooling depth. Upon hitting this floor, the only route to lowering the mode temperature still further would be the design and realization of improved low-noise transistor gates/channels, or whatever other active device structure or state of matter (beyond 2-DEGs) that can be co-opted to serve as an improved ACL, whose noise parameters can more severely reduce $T_{\text{cold}}$ as per Eq. 1.
@@
Further experimental details: Our implemented experimental set-up is shown in Fig. 4. However, when solely demonstrating mode cooling, (i) the cavity need not be loaded with an EPR sample, and (ii), the EPR bridge’s circulator can be removed with the monitoring LNA (inside the dashed green box in Fig. 4) connected directly to the critically-coupled port’s stub tuner. Our resonator is a cylindrical brass cavity, internally plated with silver, whose internal volume is partially filled by a concentrically located ring of monocrystalline sapphire (supplied by J-Crystal Photoelectric Technology, China) whose c-axis lies parallel to the cavity’s axis of rotational symmetryWu et al. (2020a). This ring supports a TE01δ mode at 1.45 GHz, whose intrinsic quality factor $Q_{\text{cav.}}$ = 164,000, providing a loaded quality factor at critical coupling ($\kappa_{\text{crit.}}=1$) of $Q_{\text{crit.}}=Q_{\text{cav.}}/2$ = 82,000. The frequency $f_{0}$ and field profile of this mode were accurately modeled using COMSOL MultiphysicsOxborrow (2007). Unlike the superheterodyne receiver (incorporating a narrowly filtered 70-MHz IF strip) used in the previous cavity-cooling works by Wu and Ng et al.Wu et al. (2021); Ng et al. (2021), we here use a simpler high-gain homodyne receiver to monitor the amplitude/excitation of the TE01δ, recording the receiver’s output into a Tektronix TBS 1102B-EDU oscilloscope.
@@
It is worth pointing out that both the cooling and the monitoring of the TE01δ mode could alternatively be implemented using just a single port attached to a single LNA, if this port’s coupling to the mode were made dynamically adjustable. The LNA would be alternatively connected (“toggled”), at sufficient speed and glitch-free, between (say) a larger and a smaller coupling loop. However, given that our cavity already had two separate coaxial ports (feedthroughs), it was simpler, with the components immediately available to us, to use both of these ports set to different fixed couplings $\kappa$, namely: (i) an over-coupled “cooling” port ($\kappa=\kappa_{\text{over}}=3.8$), connected through a low-insertion-loss single-pole double-throw (SPDT) switch (viz. a Qorvo RFSW1012 mounted on an evaluation board) to either the ACL (viz. the input of a Qorvo QPL9547 LNA mounted on an eval. board) or to an “open” termination –the choice being determined by the control voltage applied to the RFSW1012 switch; and (ii) a “monitoring/interrogation” port, critically coupled ($\kappa=\kappa_{\text{crit.}}=1$) to the TE01δ mode through a Maury Microwave 1819A triple-stub tuner, and used to monitor the mode’s field amplitude (and thus noise temperature). These two different ports are depicted, conceptually, in Fig. 2(b); their experimental embodiments are displayed in Figs. 4(a) and 4(c). The coupling factors of both ports (cooling and monitoring), as well as the quality factor of the sapphire resonator’s TE01δ mode, were accurately determined using an HP8753C vector network analyzer (VNA)Kajfez and Hwan (1984); Chua and Mirshekar-Syahkal (2003).
@@
Figure 8: (a) Zero-field (ZF) tr-EPR signal of Pc:PTP’s $T_{\text{x}}-T_{\text{z}}$ transitionWu et al. (2019) without (red) and with (blue) pre-cooling. (b) Thermal noise (in the time domain) extracted from each of these ZF tr-EPR signals and, below, their corresponding spectral densities; as before, blue and red correspond to the pre-cooled (with) and ambient (without) conditions, respectively. (c) Simplified Jablonski (= energy-level) diagram of the ZF sublevels ($T_{\text{x}}$, $T_{\text{y}}$, $T_{\text{z}}$) of pentacene’s triplet ground state, where D and E denote the ZF splitting (ZFS) parameters. (d) Modeled noise power reduction ($\Delta P$) at the homodyne receiver’s input as a function of the noise temperature of the EPR cavity’s operational mode, $T_{\text{mode}}$. Since its uncertainty is lower, the estimated depth of cooling presented here ($\Delta P_{\text{exp-Cool}}$) corresponds to that achieved by active pre-cooling alone (see Table S2 in the Supplemental Materialsup (2024)). The depth’s standard error is indicated by dashed cornflower-blue lines.
The reduction in thermal noise will persist so long as the cooling LNA is connected. Upon disconnecting it and connecting the over-coupling loop (port) instead to a reflective “open”, the mode regains its higher quality factor, namely $Q_{\text{crit.}}$, and warms up on a timescale set by the mode’s photon lifetimeAspelmeyer et al. (2014), namely $\tau_{\text{warm up}}=\frac{Q_{\text{crit.}}}{2\pi f_{0}}.$ Note that this is substantially longer than the time required for cool down. [Parenthetically: any perfectly reflecting impedance, like a “short”, could serve just at well as the open that we employ here.] Since the monitoring port is critically coupled to the TE01δ mode, this mode’s temperature will settle at the (equally weighted) average of $T_{\text{hot}}=290$ K and $T_{\text{mon.}}$, namely $T_{\text{mode}}^{\text{ambient}}=257.6$ K for our particular experiment.
@@ V Results
Preliminaries: We verified the noise characteristics of our cooling LNA (viz. a Qorvo QPL9547EVB-01 operating at room temperature) using the industry-standard Y-factor methodHewlett Packard / Agilent (1983). This included connecting the LNA’s input to a passive load, well matched to 50 $\Omega$, both at room temperature and when immersed in a small dewar filled with liquid nitrogen. Within experimental error (taking noise generated by losses along lengths of line into account), our own measurements were wholly consistent with the manufacturer’s specification sheet. When used as an amplifier, this LNA exhibits, at 1.45 GHz, an impressive noise figure of just 0.182 dB (for $\Gamma_{\text{s}}=0$; referenced on the device’s leads), equivalent to an effective noise temperature (referred to input) of just $T_{\text{eff.}}=12.4$ K; see Table S1 in the Supplemental Material sup (2024) for this LNA’s full noise parameters interpolated to 1.45 GHz, and Refs. Dunleavy (2009); Modelithics + Mini-Circuits (2015) for associated definitions. Note that $T_{\text{eff.}}$, calculated using Eq. 2, is substantially smaller that the temperature of the same device when configured as a cold load, namely $T_{\text{cold}}$ $=$ 29.1 K, calculated using Eq. 1. Since the QPL9547 MMIC was designed/optimized by its manufacturer for use as a LNA, this should not be hugely surprising.
@@
Mode Cooling: The effect of connecting and then disconnecting the over-coupled cooling LNA to our cavity’s TE01δ mode at $f_{0}=$ 1.45 GHz is shown in Fig. 6. From the difference in the time average of the measured noise voltage squared for the mode under cooled and ambient conditions, the depth of cooling is estimated to be $\Delta P_{\text{exp-Cool}}=-3.5\pm 0.4$ dB. This agrees reasonably well with both our theoretical estimate $\Delta P_{\text{model}}$ and our analysis using the noise spectrum from experimental time-resolved-EPR data, namely $\Delta P_{\text{exp-EPR}}$, as shown in Fig. 8(b). All our estimates are summarized in Table S2 of the Supplemental Materialsup (2024). Upon curve-fitting to the rise in the measured noise voltage squared as a function of time, the characteristic time quantifying the persistence of the mode’s coldness after disconnection from the cooling LNA is estimated to be $\tau_{\text{warm up}}^{\text{exp.}}=9\pm 1~{}\mu$s; see Fig. S9 in the Supplemental Materialsup (2024).
@@
To demonstrate the utility of active pre-cooling, we studied the signal-to-noise ratio (SNR) and noise spectrum of traces obtained from performing tr-EPR on a photo-excited sample of pentacene-doped para-terphenyl (Pc:PTP) at room temperature in zero applied magnetic field. This sample and the complete spectrometer around it are shown in Fig. 4. Fig. 8(a) displays tr-EPR signals for the Pc:PTP triplet’s Tx-Tz transition at 1.4495 GHz [as shown in Fig. 8(c)] both without (red) and with (blue) pre-cooling. The noise on each of the measured tr-EPR signals shown in Fig. 8(a) was extractedHassan and Anwar (2010) from the causal EPR response by first (i) generating a smoothed-out version of the measured data by boxcar averaging, the width of each averaging box being set at 100 ns, then (ii) subtracting this version from the directly measured data to provide the traces shown at the top of Fig. 8(b). Their respective spectral densities (modulus squared of Fourier transform) are displayed underneath. Here, only the section dominated by white noise, above the corner frequency of the observed $1/f$ noise (at around 1 MHz) generated by the receiver’s LNAs, is shown. These plots attest to the significant reduction in white (thermal) noise that pre-cooling can provide, the SNR (immediately after pre-cooling) being enhanced by $3.5^{+0.4}_{-0.4}$ dB. The relationship between the reduction in the noise power and the achieved noise temperature of the TE01δ mode at 1.45 GHz is shown in Fig. 8(d). In our demonstration, the mode temperature drops from an initial $T^{\text{ambient}}_{\text{mode}}~{}=~{}257.6$ K to a pre-cooled $T_{\text{mode}}^{\text{cooled}}~{}=~{}122.6$ K (measured), corresponding to a reduction in the expected number of thermal photons from 3699 to 1761.
@@
Restricting ourselves to work done on bench-tops at room temperature, we here briefly compare the performance achieved by our demonstration of active pre-cooling (APC) to that achieved by removing thermal photons via stimulated absorption in an absorptively spin-polarised medium resident inside the cavity. In the literature, this latter approach has been dubbed “masar”Wu et al. (2021) or “anti-maser”Blank et al. (2023). Using photo-excited Pc:PTP as the absorber, Wu et al.Wu et al. (2021) cooled a TE01δ mode at 1.45 GHz down to around 50 K –though only in bursts. Using NV- diamond, Ng et al.Ng et al. (2021) and Day et al.Day et al. (2024) cooled the same type of mode continuously to 188 K at 2.9 GHz and to 70 K at 9.8 GHz, respectively. Given that these temperatures straddle what we here achieve using APC, one may conclude that, currently, there is no clear winner. Both approaches can assuredly be improved upon through focused material science and device engineering. Beyond cost and convenience, a salient advantage of APC is that the disturbance on the microwave cavity imposed by the cooling mechanism can be removed at the flick of a (microwave) switch. Methods based on internal stimulated absorption are, in contrast, beholden to the particular spin dynamics of the cooling medium. One could however use an external masar/anti-maser (operated either in CW mode or quasi-continuouslyWu et al. (2020b)) as the ACL to implement pre-cooling instead of a HEMT-based LNA, with the connection between the resonator and the anti-maser’s coupling port gated by a microwave switch. As with other pulse-based techniques at microwave frequencies, a successful implementation of APC relies heavily on the availability of a sufficiently fast, low-loss and glitch-free switch.
@@
To conclude, we have demonstrated a method for actively pre-cooling the modes of a microwave cavity by temporarily over-coupling them to the input of a low-noise amplifier (LNA) serving as an active cold load (ACL). Beyond tr-EPR, other pulsed-EPR techniques such as electron spin-echo envelope modulation (ESEEM), hyperfine sublevel correlation spectroscopy (HYSCORE), and pulsed electron-electron double resonance (PELDOR) could be accelerated by APC, reducing measurement timesPoole and Farach (2019). Furthermore, our analysis identifies how an LNA that is directly coupled to a cavity (i.e., when no “protective” isolator/circulator is inserted in between), can play an active and beneficial role in reducing the number of thermal photons occupying the cavity’s modes (within the LNA’s bandwidth), so reducing measurement noise.
@@
The use of cryogenic ACLs, operating at (say) liquid-helium temperature yet providing input noise temperatures and thus mode cooling down into the mK regime, is perfectly conceivable. Unfortunately, due to self-heating within the device’s active channel/layer, the noise temperatures of what are currently the quietest semiconductor-based LNAs hit, upon physical cooling, plateaux of a few KArdizzi et al. (2022); Ardizzi (2022); Zeng et al. (2024). The construction of effective cryo-ACLs to achieve mK-level mode cooling may well require new materials and/or architectures (and the equipment to fabricate such), optimized expressly for mode-cooling –as opposed to the repurposing of existing devices (designed with other applications mind) like what our current exploration has been based upon. At a more general, sociologically level, acknowledging that the levels of cooling we have here achieved using APC are still far away from the single-photon regime, it is hoped that this paper induces greater cross-fertilization between wisdom embedded within the discipline of microwave radiometryCakaj et al. (2011); Weissbrodt (2017), activities across the EPR/NMR communitiesHyde (2005); Chen et al. (2024), and hardware innovations in quantum computing/detection, particularly those targeting room-temperature operationHenschel et al. (2010); Pezzagna and Meijer (2021).

### v3

@@ IV Switchover followed by Interrogation
The well-known Landau-Zener (L-Z) model[53, 54] provides quantitative insight into the location of this balance. Since, for the relatively high-$Q$ (by room-temperature standards) sapphire-loaded dielectric resonator that we use in our demonstration, the distances in frequency between the operational microwave mode and its neighboring ones correspond to many thousands of cavity linewidths, a switching time that is a tenth or a hundredth of the cavity lifetime easily satisfies adiabaticity according to the L-Z formula (for the probability of a non-adiabatic transition). Note also that even if adiabaticity is compromised by using a very high over-coupling factor (so perturbing the frequencies and shapes of the cavity electromagnetic modes) and/or a too-rapid switchover, the pre-cooling effect is still protected by the fact that those modes closest in frequency to the operational mode, and thus those that stand to get substantially mixed with it (at anti-crossings), will also lie within the bandwidth of the ACL, and thus stand to be cooled by it also (though admittedly their couplings to the ACL and thus their cooling depths will vary). The mixing in of these nearby modes with the operational mode should thus not substantially raise the temperature. In practice, there is a limit to the amount of switchable external coupling (usually in the form of metal stubs or loops) that can be added to a cavity without reducing the intrinsic quality factor (i.e. photon lifetime) of its modes when the external coupling is switched out (disconnected). The frequency shift $\Delta f$ associated with the imperfect switching in and out of an overcoupled ACL, is shown in Fig. 5. If the cavity’s mode structure gets excessively scrambled, more distant “hot” modes beyond the ACL’s cooling bandwidth risk getting mixed into the operational mode at switchover so raising its temperature. We point out here parenthetically that the engineering challenge of achieving a large change in coupling while incurring only a small $\Delta f$ (or at least the avoidance of anticrossings and mode scrambling) draws parallels to the realization of switchable attenuators exhibiting low dispersion, for which proven design strategies/solutions are known. Based on our crude L-Z analysis, there would appear to be at least two decades of headroom for raising $\kappa_{\text{over}}$ to the point of “diminishing returns”, where it is no longer the external coupling but the noise temperature of the room-temperature ACL itself that limits the cooling depth. Upon hitting this floor, the only route to lowering the mode temperature still further would be the design and realization of improved low-noise transistor gates/channels, or whatever other active device structure or state of matter (beyond 2-DEGs) that can be co-opted to serve as an improved ACL, whose noise parameters can more severely reduce $T_{\text{cold}}$ as per Eq. 1.
@@
Further experimental details: Our implemented experimental set-up is shown in Fig. 4. However, when solely demonstrating mode cooling, (i) the cavity need not be loaded with an EPR sample, and (ii), the EPR bridge’s circulator can be removed with the monitoring LNA (inside the dashed green box in Fig. 4) connected directly to the critically-coupled port’s stub tuner. Our resonator is a cylindrical brass cavity, internally plated with silver, whose internal volume is partially filled by a concentrically located ring of monocrystalline sapphire (supplied by J-Crystal Photoelectric Technology, China) whose c-axis lies parallel to the cavity’s axis of rotational symmetry[52]. This ring supports a TE01δ mode at 1.45 GHz, whose intrinsic quality factor $Q_{\text{cav.}}$ = 164,000, providing a loaded quality factor at critical coupling ($\kappa_{\text{crit.}}=1$) of $Q_{\text{crit.}}=Q_{\text{cav.}}/2$ = 82,000. The frequency $f_{0}$ and field profile of this mode were accurately modeled using COMSOL Multiphysics[55]. Unlike the superheterodyne receiver (incorporating a narrowly filtered 70-MHz IF strip) used in the previous cavity-cooling works by Wu and Ng et al.[1, 2], we here use a simpler high-gain homodyne receiver to monitor the amplitude/excitation of the TE01δ, recording the receiver’s output into a Tektronix TBS 1102B-EDU oscilloscope.
@@
It is worth pointing out that both the cooling and the monitoring of the TE01δ mode could alternatively be implemented using just a single port attached to a single LNA, if this port’s coupling to the mode were made dynamically adjustable. The LNA would be alternatively connected (“toggled”), at sufficient speed and glitch-free, between (say) a larger and a smaller coupling loop. However, given that our cavity already had two separate coaxial ports (feedthroughs), it was simpler, with the components immediately available to us, to use both of these ports set to different fixed couplings $\kappa$, namely: (i) an over-coupled “cooling” port ($\kappa=\kappa_{\text{over}}=3.8$), connected through a low-insertion-loss single-pole double-throw (SPDT) switch (viz. a Qorvo RFSW1012 mounted on an evaluation board) to either the ACL (viz. the input of a Qorvo QPL9547 LNA mounted on an eval. board) or to an “open” termination –the choice being determined by the control voltage applied to the RFSW1012 switch; and (ii) a “monitoring/interrogation” port, critically coupled ($\kappa=\kappa_{\text{crit.}}=1$) to the TE01δ mode through a Maury Microwave 1819A triple-stub tuner, and used to monitor the mode’s field amplitude (and thus noise temperature). These two different ports are depicted, conceptually, in Fig. 2(b); their experimental embodiments are displayed in Figs. 4(a) and 4(c). The coupling factors of both ports (cooling and monitoring), as well as the quality factor of the sapphire resonator’s TE01δ mode, were accurately determined using an HP8753C vector network analyzer (VNA)[56, 57].
@@
Figure 8: (a) Zero-field (ZF) tr-EPR signal of Pc:PTP’s $T_{\text{x}}-T_{\text{z}}$ transition[58] without (red) and with (blue) pre-cooling. (b) Thermal noise (in the time domain) extracted from each of these ZF tr-EPR signals and, below, their corresponding spectral densities; as before, blue and red correspond to the pre-cooled (with) and ambient (without) conditions, respectively. (c) Simplified Jablonski (= energy-level) diagram of the ZF sublevels ($T_{\text{x}}$, $T_{\text{y}}$, $T_{\text{z}}$) of pentacene’s triplet ground state, where D and E denote the ZF splitting (ZFS) parameters. (d) Modeled noise power reduction ($\Delta P$) at the homodyne receiver’s input as a function of the noise temperature of the EPR cavity’s operational mode, $T_{\text{mode}}$. Since its uncertainty is lower, the estimated depth of cooling presented here ($\Delta P_{\text{exp-Cool}}$) corresponds to that achieved by active pre-cooling alone (see Table S2 in the Supplemental Material[46]). The depth’s standard error is indicated by dashed cornflower-blue lines.
The reduction in thermal noise will persist so long as the cooling LNA is connected. Upon disconnecting it and connecting the over-coupling loop (port) instead to a reflective “open”, the mode regains its higher quality factor, namely $Q_{\text{crit.}}$, and warms up on a timescale set by the mode’s photon lifetime[59], namely $\tau_{\text{warm up}}=\frac{Q_{\text{crit.}}}{2\pi f_{0}}.$ Note that this is substantially longer than the time required for cool down. [Parenthetically: any perfectly reflecting impedance, like a “short”, could serve just at well as the open that we employ here.] Since the monitoring port is critically coupled to the TE01δ mode, this mode’s temperature will settle at the (equally weighted) average of $T_{\text{hot}}=290$ K and $T_{\text{mon.}}$, namely $T_{\text{mode}}^{\text{ambient}}=257.6$ K for our particular experiment.
@@ V Results
Preliminaries: We verified the noise characteristics of our cooling LNA (viz. a Qorvo QPL9547EVB-01 operating at room temperature) using the industry-standard Y-factor method[60]. This included connecting the LNA’s input to a passive load, well matched to 50 $\Omega$, both at room temperature and when immersed in a small dewar filled with liquid nitrogen. Within experimental error (taking noise generated by losses along lengths of line into account), our own measurements were wholly consistent with the manufacturer’s specification sheet. When used as an amplifier, this LNA exhibits, at 1.45 GHz, an impressive noise figure of just 0.182 dB (for $\Gamma_{\text{s}}=0$; referenced on the device’s leads), equivalent to an effective noise temperature (referred to input) of just $T_{\text{eff.}}=12.4$ K; see Table S1 in the Supplemental Material [46] for this LNA’s full noise parameters interpolated to 1.45 GHz, and Refs. [40, 41] for associated definitions. Note that $T_{\text{eff.}}$, calculated using Eq. 2, is substantially smaller that the temperature of the same device when configured as a cold load, namely $T_{\text{cold}}$ $=$ 29.1 K, calculated using Eq. 1. Since the QPL9547 MMIC was designed/optimized by its manufacturer for use as a LNA, this should not be hugely surprising.
@@
Mode Cooling: The effect of connecting and then disconnecting the over-coupled cooling LNA to our cavity’s TE01δ mode at $f_{0}=$ 1.45 GHz is shown in Fig. 6. From the difference in the time average of the measured noise voltage squared for the mode under cooled and ambient conditions, the depth of cooling is estimated to be $\Delta P_{\text{exp-Cool}}=-3.5\pm 0.4$ dB. This agrees reasonably well with both our theoretical estimate $\Delta P_{\text{model}}$ and our analysis using the noise spectrum from experimental time-resolved-EPR data, namely $\Delta P_{\text{exp-EPR}}$, as shown in Fig. 8(b). All our estimates are summarized in Table S2 of the Supplemental Material[46]. Upon curve-fitting to the rise in the measured noise voltage squared as a function of time, the characteristic time quantifying the persistence of the mode’s coldness after disconnection from the cooling LNA is estimated to be $\tau_{\text{warm up}}^{\text{exp.}}=9\pm 1~{}\mu$s; see Fig. S9 in the Supplemental Material[46].
@@
To demonstrate the utility of active pre-cooling, we studied the signal-to-noise ratio (SNR) and noise spectrum of traces obtained from performing tr-EPR on a photo-excited sample of pentacene-doped para-terphenyl (Pc:PTP) at room temperature in zero applied magnetic field. This sample and the complete spectrometer around it are shown in Fig. 4. Fig. 8(a) displays tr-EPR signals for the Pc:PTP triplet’s Tx-Tz transition at 1.4495 GHz [as shown in Fig. 8(c)] both without (red) and with (blue) pre-cooling. The noise on each of the measured tr-EPR signals shown in Fig. 8(a) was extracted[49] from the causal EPR response by first (i) generating a smoothed-out version of the measured data by boxcar averaging, the width of each averaging box being set at 100 ns, then (ii) subtracting this version from the directly measured data to provide the traces shown at the top of Fig. 8(b). Their respective spectral densities (modulus squared of Fourier transform) are displayed underneath. Here, only the section dominated by white noise, above the corner frequency of the observed $1/f$ noise (at around 1 MHz) generated by the receiver’s LNAs, is shown. These plots attest to the significant reduction in white (thermal) noise that pre-cooling can provide, the SNR (immediately after pre-cooling) being enhanced by $3.5^{+0.4}_{-0.4}$ dB. The relationship between the reduction in the noise power and the achieved noise temperature of the TE01δ mode at 1.45 GHz is shown in Fig. 8(d). In our demonstration, the mode temperature drops from an initial $T^{\text{ambient}}_{\text{mode}}~{}=~{}257.6$ K to a pre-cooled $T_{\text{mode}}^{\text{cooled}}~{}=~{}122.6$ K (measured), corresponding to a reduction in the expected number of thermal photons from 3699 to 1761.
@@
Restricting ourselves to work done on bench-tops at room temperature, we here briefly compare the performance achieved by our demonstration of active pre-cooling (APC) to that achieved by removing thermal photons via stimulated absorption in an absorptively spin-polarised medium resident inside the cavity. In the literature, this latter approach has been dubbed “masar”[1] or “anti-maser”[23]. Using photo-excited Pc:PTP as the absorber, Wu et al.[1] cooled a TE01δ mode at 1.45 GHz down to around 50 K –though only in bursts. Using NV- diamond, Ng et al.[2] and Day et al.[5] cooled the same type of mode continuously to 188 K at 2.9 GHz and to 70 K at 9.8 GHz, respectively. Given that these temperatures straddle what we here achieve using APC, one may conclude that, currently, there is no clear winner. Both approaches can assuredly be improved upon through focused material science and device engineering. Beyond cost and convenience, a salient advantage of APC is that the disturbance on the microwave cavity imposed by the cooling mechanism can be removed at the flick of a (microwave) switch. Methods based on internal stimulated absorption are, in contrast, beholden to the particular spin dynamics of the cooling medium. One could however use an external masar/anti-maser (operated either in CW mode or quasi-continuously[62]) as the ACL to implement pre-cooling instead of a HEMT-based LNA, with the connection between the resonator and the anti-maser’s coupling port gated by a microwave switch. As with other pulse-based techniques at microwave frequencies, a successful implementation of APC relies heavily on the availability of a sufficiently fast, low-loss and glitch-free switch.
@@
To conclude, we have demonstrated a method for actively pre-cooling the modes of a microwave cavity by temporarily over-coupling them to the input of a low-noise amplifier (LNA) serving as an active cold load (ACL). Beyond tr-EPR, other pulsed-EPR techniques such as electron spin-echo envelope modulation (ESEEM), hyperfine sublevel correlation spectroscopy (HYSCORE), and pulsed electron-electron double resonance (PELDOR) could be accelerated by APC, reducing measurement times[18]. Furthermore, our analysis identifies how an LNA that is directly coupled to a cavity (i.e., when no “protective” isolator/circulator is inserted in between), can play an active and beneficial role in reducing the number of thermal photons occupying the cavity’s modes (within the LNA’s bandwidth), so reducing measurement noise.
@@
The use of cryogenic ACLs, operating at (say) liquid-helium temperature yet providing input noise temperatures and thus mode cooling down into the mK regime, is perfectly conceivable. Unfortunately, due to self-heating within the device’s active channel/layer, the noise temperatures of what are currently the quietest semiconductor-based LNAs hit, upon physical cooling, plateaux of a few K[63, 64, 65]. The construction of effective cryo-ACLs to achieve mK-level mode cooling may well require new materials and/or architectures (and the equipment to fabricate such), optimized expressly for mode-cooling –as opposed to the repurposing of existing devices (designed with other applications mind) like what our current exploration has been based upon. At a more general, sociologically level, acknowledging that the levels of cooling we have here achieved using APC are still far away from the single-photon regime, it is hoped that this paper induces greater cross-fertilization between wisdom embedded within the discipline of microwave radiometry[66, 31], activities across the EPR/NMR communities[67, 68], and hardware innovations in quantum computing/detection, particularly those targeting room-temperature operation[69, 70].